# Initialization Schemes for Kolmogorov–Arnold Networks: An Empirical Study

### v1

@@ 1 Introduction
Kolmogorov–Arnold Networks (KANs) (Liu et al., 2025) have recently emerged as an alternative backbone architecture to Multilayer Perceptrons (MLPs), drawing inspiration from the Kolmogorov–Arnold representation theorem (Kolmogorov, 1957) in a manner analogous to how the learning of MLPs relies on universal approximation theorems. Unlike MLPs, which use fixed nonlinear activation functions and trainable synaptic weights, KANs comprise grid-dependent trainable activation functions. This provides them with flexibility in modeling complex nonlinear relationships, while requiring fewer and smaller layers. Since their introduction, KANs have found numerous applications, often surpassing the performance of their MLP-based counterparts (Yu et al., 2024; Poeta et al., 2024). There have been many notable results in scientific problem-solving domains, including function fitting and symbolic regression (Liu et al., 2024; Shukla et al., 2024), partial differential equations (PDEs) (Shukla et al., 2024; Rigas et al., 2024; Wang et al., 2025a) and operator learning (Abueidda et al., 2025; Shukla et al., 2024), among other applications (Howard et al., 2024; Kundu et al., 2024; Kashefi, 2025).
@@
Beyond these benchmarks, there has also been significant progress in the theoretical understanding of KANs (Zhang and Zhou, 2025; Alter et al., 2025; Wang et al., 2025b). However, one important theoretical and practical aspect that remains understudied pertains to their initialization strategies. Current literature mainly relies on the standard initialization method proposed in the introductory KAN paper (Liu et al., 2025), or explores alternative KAN variants such as Chebyshev-based formulations (Rigas et al., 2025). This highlights a clear gap and motivates an investigation into more effective initialization approaches for the standard spline-based architecture. Effective initialization is crucial, as a good “initial guess” for the network weights can significantly accelerate training (Mishkin and Matas, 2016; Skorski et al., 2021) and prevent early saturation of hidden layers (Glorot and Bengio, 2010). However, despite extensive research into initialization methods for MLP-based architectures, these results cannot be directly applied to KANs. Furthermore, even within MLP-based architectures, initialization methods often require separate consideration depending on the specific architecture design (Huang et al., 2020), activation function (He et al., 2015), or even on a complete case-by-case basis (Skorski et al., 2021).
@@
In response to this research gap, this work explores initialization strategies for the standard, spline-based KAN architecture. Drawing parallels with MLPs, we propose variance-preserving schemes inspired by LeCun (LeCun et al., 1998) and Glorot (Glorot and Bengio, 2010) initializations, including a variant that employs batch-normalized (Ioffe and Szegedy, 2015) spline basis functions. In addition, recognizing that theoretical frameworks may not always align with empirical performance (Mishkin and Matas, 2016), we further propose an empirical family of power-law initializations with tunable exponents. To evaluate these approaches, we conduct extensive grid searches on function fitting tasks and forward PDE problems. Beyond final error metrics, we further compare the different initialization schemes through an analysis of training dynamics, examining the evolution of loss curves and the Neural Tangent Kernel (NTK) spectrum Jacot et al. (2018); Wang et al. (2022) for representative architectures. Finally, we evaluate said architectures on a subset of the Feynman dataset (Udrescu and Tegmark, 2020), which, although widely used for symbolic regression, is formulated here as a function fitting benchmark as in Liu et al. (2025).
@@ 2.2 Related Work
In the existing KAN literature, initialization strategies have only been explored in certain KAN variants (see, e.g., Guilhoto and Perdikaris (2025)), while the standard spline-based architecture has not yet received dedicated attention in this regard. A natural starting point for studying initialization is to follow the historical developments in MLP-based architectures, beginning with variance-preserving schemes such as those proposed by LeCun (LeCun et al., 1998) and Glorot (Glorot and Bengio, 2010), which stabilize activation variance across layers and mitigate progressive vanishing or explosion. Alternatively, architectural modifications such as batch normalization (Ioffe and Szegedy, 2015) have been introduced to maintain stable activation distributions and reduce sensitivity to initialization. Within the KAN family, Glorot-inspired initialization has been applied successfully to Chebyshev-based variants (Rigas et al., 2025), though this setting differs substantially from the spline-based case studied here, since it removes the residual term of Eq. (1) and employs a completely different basis function. Consequently, it remains unclear whether such strategies directly transfer to the standard KAN formulation, motivating the investigation presented in this work. To the best of our knowledge, the present work provides the first systematic study of initialization strategies for spline-based KANs.
@@ 3.1 Proposed Initializations
If we further assume that each component of $\mathbf{x}$ is drawn from a given distribution (e.g., the uniform distribution $\mathcal{U}\left(-1,1\right)$, as is often the case in tasks like function fitting or PDE solving), then all statistical quantities in Eq. (2) can be evaluated directly, except for $\mu_{B}^{(0)}$. Due to the dependence of the spline-basis functions on the underlying grid, no general analytic expression exists for $\sigma_{b}$. This leads to two practical alternatives: one may either estimate $\mu_{B}^{(0)}$ numerically by sampling a large number of input points from the assumed distribution at initialization, or set the expectation value to unity by modifying the architecture of the KAN layer to use batch-normalized spline basis functions, defined as
@@
where the expectation values are computed over the current batch during each forward pass. We will refer to the former alternative as “LeCun–numerical” initialization, while the latter is referred to as “LeCun–normalized” initialization.
@@
where $\alpha$ and $\beta$ are tunable exponents selected from the set $\left\{0.0,0.25,\dots,1.75,2.0\right\}$. The motivation behind this empirical scheme is to perform a grid search over $(\alpha,\beta)$ configurations in order to identify trends or specific exponent pairs that consistently improve training speed and convergence, potentially revealing an effective heuristic for initializing KANs.
@@ 3.2 Experimental Setup
We evaluate initialization strategies on two benchmark families: function fitting tasks and forward PDE problems. For function fitting, we use five two-dimensional target functions and train for 2,000 epochs, while for PDEs we consider the Allen–Cahn equation, Burgers’ equation, and the two-dimensional Helmholtz equation, using KANs trained for 5,000 epochs within the Physics-Informed Machine Learning (PIML) framework (Raissi et al., 2019). Across both benchmarks, performance is measured using the final training loss and the relative $L^{2}$ error with respect to the reference solution. For the purposes of the initial grid search, we test architectures with 1–4 hidden layers, widths equal to $2^{i}$ for $i=1,\dots,6$, and grid sizes $G\in\{5,10,20,40\}$ for function fitting, while for PDEs we restrict to $G\in\{5,10,20\}$. All experiments presented herein are repeated with five random seeds, except in the power-law grid search where we use three seeds to reduce computational cost, and we report the median outcome across runs. Further implementation details, including the explicit formulas of the target functions and the PDE setups, are provided in Appendix C. All experiments are implemented in JAX (Bradbury et al., 2018), with KANs trained using the jaxKAN framework (Rigas and Papachristou, 2025). Training is performed on a single NVIDIA GeForce RTX 4090 GPU.
@@ 4.1 Grid-Search Results
We observe that, in nearly all cases, a suitable power-law initialization outperforms the baseline scheme across both metrics. In contrast, the LeCun-inspired strategies yield more mixed outcomes: while they rarely outperform the baseline on small architectures, the normalized variant in particular becomes more effective as depth, width and grid size increase, in some cases providing improvements of over two orders of magnitude. The Glorot-inspired initialization performs consistently better than both the baseline and the LeCun-based schemes, with success rates above 60%–75% across all functions, though it still falls short of the power-law family. Importantly, the more favorable power-law configurations tend to correspond to relatively small values of $\alpha$ (e.g., near 0.25) combined with larger $\beta$ values (i.e., $\geq 1.5$), leading to substantially lower final losses. The complete grid-search results, where these trends can be directly observed, are provided in Appendix D and the supplementary material.
For the PDE benchmarks, the same procedure produced 56,882 trained models, which reduced to 18,360 representative results after aggregation and 1,080 final entries after selecting the best $(\alpha,\beta)$ per setting. Table 2 summarizes the outcomes in terms of final training loss, relative $L^{2}$ error, and their joint improvement over the baseline initialization. The power-law initialization dominates in nearly all cases when judged by final loss and also achieves strong results when both metrics are considered simultaneously, although the improvements are less widespread than in the function fitting benchmarks. This is expected, since favorable initialization for function fitting does not directly transfer to PDE problems, where derivatives are involved and introduce additional complexity (Wang et al., 2024). The Glorot-inspired scheme provides a competitive alternative, with success rates of roughly 50–75% depending on the PDE, while both LeCun-based variants prove largely ineffective, with the normalized version in particular failing to improve over the baseline at all. Interestingly, this stands in contrast to the function fitting benchmarks, where the normalized variant outperformed the numerical one, again proving that PDE problems introduce qualitatively different dynamics in how initialization interacts with training. Finally, while the specific $(\alpha,\beta)$ configurations that perform best in PDEs differ from those in function fitting, the general trend persists: smaller values of $\alpha$ and larger values of $\beta$ are favored. The complete grid-search results supporting these observations are provided in Appendix D and the supplementary material.
@@ 4.2 Training Dynamics Analysis
The grid-search experiments established that Glorot- and power-law-based schemes provide strong alternatives to the baseline initialization, with the latter yielding the most consistent improvements overall. To better understand the mechanisms behind these trends, we next examine training dynamics in greater detail, focusing only on the baseline, Glorot-inspired, and power-law schemes. We begin with the function fitting benchmarks: Figure 1 shows the evolution of the training loss for two representative settings, a “small” architecture ($G=5$, two hidden layers with 8 neurons each) and a “large” architecture ($G=20$, three hidden layers with 32 neurons each). For consistency across benchmarks, we fix the power-law parameters to $\alpha=0.25$ and $\beta=1.75$, which lie within the range identified as favorable in the grid search. The curves are averaged over five seeds, with shaded regions indicating the standard error.
@@
Figure 1: Training loss curves for function fitting benchmarks under baseline, Glorot and power-law initializations. Results are averaged over five seeds, with shaded regions indicating the standard error. Top row: “small” architecture ($G=5$, two hidden layers with 8 neurons each). Bottom row: “large” architecture ($G=20$, three hidden layers with 32 neurons each).
For the small architecture, Glorot and the baseline perform similarly, while the power-law scheme clearly outperforms both by converging faster and reaching lower final losses. In contrast, for the larger architecture, the differences become much more pronounced. The baseline initialization always plateaus early, struggling to further decrease the loss, whereas Glorot maintains steady descent and power-law achieves significantly lower final losses across all functions. Importantly, the power-law initialization not only reaches the lowest losses but also does so with the steepest initial decrease, indicating faster optimization from the very first iterations. The more pronounced oscillations observed in its curves stem from the fact that the model already reaches very low loss values under a fixed learning rate, leaving no smoothing effect from adaptive adjustments.
@@
Figure 2: Training loss curves for forward PDE benchmarks under baseline, Glorot, and power-law initializations. Results are averaged over five seeds, with shaded regions indicating the standard error. Top row: “small” architecture ($G=5$, two hidden layers with 8 neurons each). Bottom row: “large” architecture ($G=20$, three hidden layers with 32 neurons each).
We repeat the same analysis for the PDE benchmarks, with results shown in Figure 2. For the smaller architecture, the three initialization schemes perform comparably, though in the Allen–Cahn case the power-law scheme achieves noticeably faster convergence, with lower variance. By contrast, in the larger architecture, the differences are striking: the baseline often stagnates at high loss values, while Glorot initialization provides steady improvement and power-law achieves the lowest overall errors, albeit with oscillations similar to those observed in the function fitting tasks. Overall, the PDE results reinforce the earlier findings, confirming that both Glorot and power-law yield significant gains over the baseline scheme, with the advantage becoming most apparent in parameter-rich models.
@@
To further probe the mechanisms underlying these differences, we complement the loss curve analysis with a study of the NTK dynamics. Since the discrepancies between initialization strategies are most pronounced in larger models, we focus here on the “large” architecture. Figure 3 shows the NTK eigenvalue spectra at initialization, at intermediate training iterations, and at convergence, for function fitting task $f_{3}(x,y)$ (results for other targets are consistent and reported in Appendix E.2). The baseline initialization yields a spectrum concentrated at large values that progressively collapses during training, suggesting poor conditioning and limited effective rank. By contrast, both the Glorot- and power-law-based schemes produce spectra that stabilize quickly and remain stable throughout training.
@@
Figure 3: Eigenvalue spectra of the NTK matrix at initialization (solid blue), intermediate iterations (dashed teal), and final iteration (dashed green) for function fitting benchmark $f_{3}(x,y)$ under different initialization strategies. Results correspond to the “large” architecture ($G=20$, three hidden layers with 32 neurons each).
@@
We also extend the NTK analysis to PDE benchmarks, focusing on the Allen–Cahn equation as a representative case (results for Burgers and Helmholtz are provided in Appendix E.2). To this end, we adopt the NTK formalism developed for PIML (Wang et al., 2022) and adapt it to account for Residual-Based Attention (RBA) weights (Anagnostopoulos et al., 2024), which are applied in the loss functions studied herein (see Appendix C.2 for details). The resulting kernel is identical to the standard PINN NTK, except that it incorporates the corresponding RBA weights (see Appendix E.1 for the full derivation). Figure 4 shows the NTK eigenvalue spectra separately for the PDE residual term (top row) and the boundary/initial conditions (bottom row).
@@
Figure 4: NTK eigenvalue spectra for the Allen–Cahn PDE benchmark under baseline, Glorot, and power-law initializations. Top row: spectra corresponding to the PDE residual term. Bottom row: spectra for the boundary/initial condition terms. Solid lines show the initialization, dashed lines show intermediate iterations, and dotted lines show the final iteration. Results correspond to the “large” architecture ($G=20$, three hidden layers with 32 neurons each).
@@
The PDE residual spectra largely mirror the function fitting case: baseline initialization yields poorly conditioned eigenvalues that collapse over training, while both Glorot and power-law maintain stability. The key difference lies in the boundary/initial condition terms, where Glorot shows some irregularities, though far less severe than the baseline. The power-law scheme stands out as the most consistent, providing stable and well-structured spectra across both PDE residual- and boundary/initial-term components.

### v2

@@ 1 Introduction
Kolmogorov–Arnold Networks (KANs) (Liu et al., 2025) have recently emerged as an alternative backbone architecture to Multilayer Perceptrons (MLPs), drawing inspiration from the Kolmogorov–Arnold representation theorem (Kolmogorov, 1957) in a manner analogous to how the learning of MLPs relies on universal approximation theorems. Unlike MLPs, which use fixed nonlinear activation functions and trainable synaptic weights, KANs comprise grid-dependent trainable activation functions. This provides them with flexibility in modeling complex nonlinear relationships, while requiring fewer and smaller layers. Since their introduction, KANs have found numerous applications, often surpassing the performance of their MLP-based counterparts (Yu et al., 2024; Poeta et al., 2024). There have been many notable results in scientific problem-solving domains, including function fitting and symbolic regression (Liu et al., 2024; Shukla et al., 2024), partial differential equations (PDEs) (Shukla et al., 2024; Rigas et al., 2024; Wang et al., 2025b) and operator learning (Abueidda et al., 2025; Shukla et al., 2024; Lee et al., 2025), among other applications (Howard et al., 2024; Kundu et al., 2024; Kashefi, 2025).
@@
Beyond these benchmarks, there has also been significant progress in the theoretical understanding of KANs (Zhang and Zhou, 2025; Alter et al., 2025; Wang et al., 2025a). However, one important theoretical and practical aspect that remains understudied pertains to their initialization strategies. Current literature mainly relies on the standard initialization method proposed in the introductory KAN paper (Liu et al., 2025), or explores alternative KAN variants such as Chebyshev-based formulations (Rigas et al., 2026). This highlights a clear gap and motivates an investigation into more effective initialization approaches for the standard spline-based architecture. Effective initialization is crucial, as a good “initial guess” for the network weights can significantly accelerate training (Mishkin and Matas, 2016; Skorski et al., 2021) and prevent early saturation of hidden layers (Glorot and Bengio, 2010). However, despite extensive research into initialization methods for MLP-based architectures, these results cannot be directly applied to KANs. Furthermore, even within MLP-based architectures, initialization methods often require separate consideration depending on the specific architecture design (Huang et al., 2020), activation function (He et al., 2015), or even on a complete case-by-case basis (Skorski et al., 2021).
@@
In response to this research gap, this work explores initialization strategies for the standard, spline-based KAN architecture. Drawing parallels with MLPs, we propose variance-preserving schemes inspired by LeCun (LeCun et al., 1998) and Glorot (Glorot and Bengio, 2010) initializations, including a variant that employs normalized spline basis functions. In addition, recognizing that theoretical frameworks may not always align with empirical performance (Mishkin and Matas, 2016), we further propose an empirical family of power-law initializations parameterized by two exponents. We perform grid searches to identify suitable exponent choices for the power-law method on function fitting and forward PDE benchmarks and then evaluate all initialization schemes on these tasks. We subsequently fix the exponents to values that lie within the identified well-performing range and select two representative architectures (in terms of parameter count) to analyze training dynamics through the evolution of training loss curves and the Neural Tangent Kernel (NTK) spectrum (Jacot et al., 2018; Wang et al., 2022). Finally, we evaluate said architectures on a subset of the Feynman dataset (Udrescu and Tegmark, 2020), which, although widely used for symbolic regression, is formulated here as a function fitting benchmark as in Liu et al. (2025).
@@ 2.2 Related Work
In the existing KAN literature, initialization strategies have only been explored in certain KAN variants (see, e.g., Guilhoto and Perdikaris (2025)), while the standard spline-based architecture has not yet received dedicated attention in this regard. A natural starting point for studying initialization is to follow the historical developments in MLP-based architectures, beginning with variance-preserving schemes such as those proposed by LeCun (LeCun et al., 1998) and Glorot (Glorot and Bengio, 2010), which stabilize activation variance across layers and mitigate progressive vanishing or explosion. Within the KAN family, Glorot-inspired initialization has been applied successfully to Chebyshev-based variants (Rigas et al., 2026), though this setting differs substantially from the spline-based case studied here, since it removes the residual term of Eq. (1) and employs a completely different basis function. Consequently, it remains unclear whether such strategies directly transfer to the standard KAN formulation, motivating the investigation presented in this work. To the best of our knowledge, the present work provides the first systematic study of initialization strategies for spline-based KANs.
@@ 3.1 Proposed Initializations
If we further assume that each component of $\mathbf{x}$ is drawn from a given distribution (e.g., the uniform distribution $\mathcal{U}\left(-1,1\right)$, as is often the case in tasks like function fitting or PDE solving), then all statistical quantities in Eq. (2) can be evaluated directly, except for $\mu_{B}^{(0)}$. Due to the dependence of the spline-basis functions on the underlying grid, no general analytic expression exists for $\sigma_{b}$. This leads to two practical alternatives: one may either estimate $\mu_{B}^{(0)}$ numerically by sampling a large number of input points from the assumed distribution at initialization, or set the expectation value to unity by modifying the architecture of the KAN layer to use normalized spline basis functions, defined as
@@
where the expectation values are computed over the layer inputs during the forward pass. We will refer to the former alternative as “LeCun–numerical” initialization, while the latter is referred to as “LeCun–normalized” initialization.
@@
where $\alpha$ and $\beta$ are tunable exponents selected from the set $\left\{0.0,0.25,\dots,1.75,2.0\right\}$. The motivation behind this empirical scheme is to perform a grid search over $(\alpha,\beta)$ configurations in order to identify trends or specific exponent pairs that consistently improve training speed and convergence. Such searches can be carried out on a per-domain basis (e.g., function fitting, forward PDEs), after which the resulting well-performing ranges may serve as reusable heuristics for future problems of the same type.
@@ 3.2 Experimental Setup
We evaluate initialization strategies on two benchmark families: function fitting tasks and forward PDE problems. For function fitting, we use five two-dimensional target functions and train for 2,000 iterations (epochs), while for PDEs we consider the Allen–Cahn equation, Burgers’ equation, and the two-dimensional Helmholtz equation, using KANs trained for 5,000 iterations within the Physics-Informed Machine Learning (PIML) framework (Raissi et al., 2019). Across both benchmarks, performance is measured using the final training loss and the relative $L^{2}$ error with respect to the reference solution. For the purposes of the initial grid search, we test architectures with 1–4 hidden layers, widths equal to $2^{i}$ for $i=1,\dots,6$, and grid sizes $G\in\{5,10,20,40\}$ for function fitting, while for PDEs we restrict to $G\in\{5,10,20\}$. All experiments presented herein are repeated with five random seeds, except in the power-law grid search where we use three seeds to reduce computational cost, and we report the median outcome across runs. Further implementation details, including the explicit formulas of the target functions and the PDE setups, are provided in Appendix C. All experiments are implemented in JAX (Bradbury et al., 2018), with KANs trained using the jaxKAN framework (Rigas and Papachristou, 2025). Training is performed on a single NVIDIA GeForce RTX 4090 GPU.
@@ 4.1 Grid-Search Results
The LeCun-inspired schemes rarely outperform the baseline on the smaller architectures, but their effectiveness increases with depth, width and grid size. In some of the larger settings, the normalized variant in particular achieves improvements exceeding two orders of magnitude relative to the baseline. However, in terms of absolute frequency, Table 1 clearly shows that the baseline still outperforms both LeCun-based variants: for more than 70% of configurations, the resulting relative $L^{2}$ error under LeCun initialization is higher than under the baseline scheme. Between the two LeCun variants, the normalized version consistently performs better than the numerical one, which is consistent with its design, as variance preservation is enforced by construction. On the other hand, the Glorot-inspired initialization performs more robustly. Across all five functions, it yields success rates of approximately 60–75% for both loss and relative $L^{2}$ error, indicating that simultaneously balancing forward- and backward-pass variances is considerably more effective than forward-variance preservation alone. The few cases where the baseline performs better occur predominantly for the smaller architectures.
As far as the power-law initialization is concerned, it exhibits the strongest overall performance. Table 1 shows that, for nearly every architecture and target function, there exists at least one $(\alpha,\beta)$ pair that outperforms the baseline, often by a substantial margin. The most favorable region is concentrated around small residual exponents, i.e., $\alpha\in\{0.25,0.5\}$, combined with larger basis-function exponents, namely $\beta\geq 1.0$. Full grid-search results illustrating these trends are provided in Appendix D and the supplementary material. Notably, even when fixing a single configuration within this region, the method remains highly robust. For example, with $(\alpha,\beta)=(0.25,1.0)$, the initialization yields simultaneous improvements in both loss and $L^{2}$ error over the baseline in 96.88% of runs for $f_{1}(x,y)$, 95.83% for $f_{2}(x,y)$, 97.92% for $f_{3}(x,y)$, 87.50% for $f_{4}(x,y)$, and 89.58% for $f_{5}(x,y)$. This indicates that once a suitable exponent range is identified for a given problem type, a fixed choice within that region can systematically outperform both the baseline and the Glorot initialization.
Following the same procedure for the PDE benchmarks, we trained 56,882 models, a number which reduced to 18,360 representative results after aggregation and 1,080 final entries after selecting the best $(\alpha,\beta)$ per setting. Table 2 summarizes the outcomes in terms of final training loss, relative $L^{2}$ error, and their joint improvement over the baseline initialization.
@@
When comparing the two LeCun-based schemes to each other, the observed behavior is essentially the inverse of what we observed in function fitting. The normalized variant, which performed better than the numerical one for function fitting, fails almost entirely in PDE problems: in nearly all configurations, it produces no improvement over the baseline, with success rates effectively equal to zero. This discrepancy can be attributed to the fact that PDE losses involve not only the network output but also its derivatives and nonlinear combinations thereof. Normalizing the spline basis propagates all multiplicative constants (the standard deviation in the denominator of Eq. (4) in this case) into all derivatives, altering the stiffness of the residuals. While the numerical variant avoids this issue, the results are still poor: although it occasionally outperforms the baseline on larger architectures, it remains generally ineffective. The Glorot-inspired initialization again shows more consistent improvements. As in function fitting, it performs significantly better than the baseline on parameter-rich architectures and the cases where it underperforms correspond to smaller models.
The power-law initialization remains the strongest of all approaches, though the advantage is less pronounced in the PDE case. For more than 70% of configurations (up to 94.44% in the Allen–Cahn equation) there exists at least one $(\alpha,\beta)$ pair that outperforms the baseline simultaneously in both loss and relative $L^{2}$ error. The preferred exponent region differs slightly from the function fitting case: while small residual exponents $\alpha$ remain strongly favored, in this case lower values of $\beta$ (typically $0.75\leq\beta\leq 1.25$) yield the best results. The complete grid-search results supporting these observations are provided in Appendix D and the supplementary material. For the previously discussed configuration $(\alpha,\beta)=(0.25,1.0)$, the power-law initialization outperforms the baseline on both metrics in 83.33% of runs for Allen–Cahn, 54.93% for Burgers’ and 59.72% for Helmholtz, with even higher success rates when each metric is considered individually.
@@ 4.2 Training Dynamics Analysis
The previous experiments established that the Glorot- and power-law-based schemes provide strong alternatives to the baseline initialization, with the latter yielding the most consistent improvements overall. In contrast, the LeCun-based variants exhibited substantially weaker performance. To better understand the mechanisms behind these trends, we next examine training dynamics in greater detail for each initialization scheme, excluding only the LeCun-normalized variant due to its complete breakdown in PDE problems for the aforementioned reasons. We begin with the function fitting benchmarks: Figure 1 shows the evolution of the training loss for two representative settings, a “small” architecture ($G=5$, two hidden layers with 8 neurons each) and a “large” architecture ($G=20$, three hidden layers with 32 neurons each). For consistency across experiments, we fix the power-law parameters to $\alpha=0.25$ and $\beta=1.75$, which lie within the range identified as favorable in the grid search for function fitting. The curves are averaged over five seeds, with shaded regions indicating the standard error.
@@
Figure 1: Training loss curves for function fitting benchmarks under baseline, LeCun-numerical, Glorot and power-law ($\alpha=0.25,\beta=1.75$) initializations. Results are averaged over five seeds, with shaded regions indicating the standard error. Top row: “small” architecture ($G=5$, two hidden layers with 8 neurons each). Bottom row: “large” architecture ($G=20$, three hidden layers with 32 neurons each).
Across all settings, the loss curves in Figure 1 reinforce the conclusions drawn from the previous experiments: the power-law initialization consistently outperforms all other schemes, converging both faster and to lower final losses. For the small architecture, the baseline and Glorot initializations behave similarly, whereas for the larger architecture Glorot has a clear advantage over the baseline. The LeCun initialization, while stable and not prone to divergence, persistently underperforms the remaining schemes in both architectures. The oscillatory behavior observed for Glorot and power-law in the large-architecture setting is a consequence of using a fixed learning rate throughout training. We intentionally avoided learning-rate scheduling because initialization and learning-rate adaptability are known to interact in subtle ways (e.g., (Yang et al., 2021)), and our goal was to isolate the effect of initialization alone. In Appendix E we provide results for the same experiments using a learning-rate scheduler; the resulting curves are significantly smoother, however the relative performance ranking of the initialization schemes remains unchanged.
The corresponding analysis for the PDE benchmarks is shown in Figure 2, where again we use the power-law parameters $\alpha=0.25$ and $\beta=1.75$, despite the grid search identifying a different optimal region for $\beta$, in order to demonstrate that the method remains robust even when not tuned specifically for PDEs. For the small architecture, all initialization schemes ultimately perform comparably, albeit at different rates. Notably, the power-law initialization exhibits markedly lower variance across seeds and, in the Allen–Cahn case in particular, reaches a minimum significantly faster than the alternatives. In the larger architecture, however, the trends closely mirror those observed in the function fitting experiments: the LeCun initialization lags behind by several orders of magnitude, the baseline also underperforms, and both Glorot and power-law provide substantial improvements. These two consistently achieve the lowest losses, with the power-law scheme displaying an advantage in the Allen–Cahn and Burgers equations. As with the function fitting results, the oscillatory behavior observed for Glorot and power-law arises from the fixed learning rate used during training. The corresponding smoothed curves obtained with learning-rate scheduling are provided in Appendix E.
@@
Figure 2: Training loss curves for forward PDE benchmarks under baseline, LeCun-numerical, Glorot, and power-law ($\alpha=0.25,\beta=1.75$) initializations. Results are averaged over five seeds, with shaded regions indicating the standard error. Top row: “small” architecture ($G=5$, two hidden layers with 8 neurons each). Bottom row: “large” architecture ($G=20$, three hidden layers with 32 neurons each).
@@
To gain further insight into why the baseline and LeCun initializations underperform while Glorot and power-law consistently succeed, we complement the loss-curve analysis with a study of NTK dynamics. Since the discrepancies between initialization strategies are most pronounced in larger models, we focus here on the “large” architecture. For the power-law method we keep $\alpha=0.25$ and $\beta=1.75$, however we also provide the spectra for other configurations in Appendix F.2. Figure 3 shows the NTK eigenvalue spectra at initialization, at intermediate training iterations, and at convergence, for the task of fitting $f_{3}(x,y)$ (results for other targets are provided in Appendix F.3).
@@
Figure 3: Eigenvalue spectra of the NTK matrix at initialization (solid blue), intermediate iterations (dashed teal), and final iteration (dashed green) for function fitting benchmark $f_{3}(x,y)$ under different initialization strategies. Results correspond to the “large” architecture ($G=20$, three hidden layers with 32 neurons each). The power-law initialization uses $\alpha=0.25,\beta=1.75$.
The spectra reveal several notable differences. The baseline initialization exhibits a spectrum dominated by very large leading eigenvalues and a steep decay, which collapses further during training, indicating poor conditioning and a low effective rank. LeCun behaves similarly but with even more extreme magnitudes. In contrast, the Glorot initialization produces a well-spread spectrum with stable leading and trailing eigenvalues throughout optimization. The power-law scheme yields an even better spectrum, closely following a power-law decay at initialization and remaining perfectly stable during training, suggesting balanced sensitivity across modes. These observations align with the performance trends reported so far: Glorot and power-law initializations induce stable, well-conditioned NTK spectra and therefore correspond to faster optimization and lower final error, while the baseline and LeCun produce highly skewed or collapsing spectra and thus consistently underperform during training.
@@
We also extend the NTK analysis to PDE benchmarks, focusing on the Allen–Cahn equation as a representative case (results for Burgers and Helmholtz are provided in Appendix F.3). To this end, we adopt the NTK formalism developed for PIML (Wang et al., 2022) and adapt it to account for Residual-Based Attention (RBA) weights (Anagnostopoulos et al., 2024), which are applied in the loss functions studied herein (see Appendix C.2 for details). The resulting kernel is identical to the standard PINN NTK, except that it incorporates the corresponding RBA weights (see Appendix F.1 for the full derivation). Figure 4 shows the NTK eigenvalue spectra separately for the PDE residual term (top row) and the boundary/initial conditions (bottom row).
@@
Figure 4: NTK eigenvalue spectra for the Allen–Cahn PDE benchmark under baseline, LeCun-numerical, Glorot, and power-law ($\alpha=0.25,\beta=1.75$) initializations. Top row: spectra corresponding to the PDE residual term. Bottom row: spectra for the boundary/initial condition terms. Solid blue lines show the initialization, dashed teal lines show intermediate iterations, and dashed green lines show the final iteration. Results correspond to the “large” architecture ($G=20$, three hidden layers with 32 neurons each).
@@
The PDE residual spectra largely mirror the function fitting case: baseline and LeCun initializations yield poorly conditioned eigenvalues that collapse over training, while both Glorot and power-law maintain stability. The key difference lies in the boundary/initial condition terms, where Glorot shows some irregularities, though far less severe than the baseline and LeCun. The power-law scheme stands out as the most consistent, providing stable and well-structured spectra across both PDE residual- and boundary/initial-term.